# Variational Mode-Driven Graph Convolutional Network for Spatiotemporal Traffic Forecasting

### v1

@@ IV-D Ablative Study
The performance of the framework has been evaluated on various combinations of decomposition hyper-parameters such as $\alpha$, and $\epsilon$, for the SD region. TABLE III details different performance metrics for horizons $3$, $6$, and $12$, and the average. It has been observed that for $\alpha=1000$, the reconstruction loss is relatively lower compared to $\alpha=2000$. Changing the $\epsilon$ value does not significantly affect the reconstruction loss, indicating that the convergence is sensitive to initial conditions, which may lead to local minima instead of global minimum. For VMGCN-v2, it is noticed that decreasing the bandwidth constraint and increasing the tolerance value improves the performance metrics. The overall cumulative effect on the reconstruction loss term in the region remains unaffected by changes in the $\epsilon$ term. However, the difference in the reconstructed loss term of modes defined as
$\frac{1}{L}\lvert\sum_{i}u(t)_{i,\epsilon 1}-\sum_{i}u(t)_{i,\epsilon 2}\rvert,$ (14)
@@
where $L$ is the total length of the signal. For this comparison, we used $\epsilon_{1}=1\times 10^{-7}$ and $\epsilon_{2}=1\times 10^{-6}$. With a bandwidth constraint of $\alpha=1000$, the loss term is measured as $1.134\times 10^{-5}$, indicating a convergence difference between these sets of decomposition parameters, with different local minima being reached. The framework performs better with $\epsilon_{2}$ compared to $\epsilon_{1}$. If the bandwidth constraint is set at $\alpha=2000$, the $L_{1}$ loss term of the mode approaches zero, indicating that at a larger bandwidth value, the role of tolerance in convergence is less significant. From the table, we may also infer that changing the bandwidth constraint has a dominant impact on the performance of the network. By eliminating the high-frequency (IMF13) in VMGCN-v1, the DNN is trained and evaluates the performance of the model. It appears that the overall performance of the network is enhanced, specifically the long-term predictions. The high-frequency component adversely affects the long-term prediction capabilities of the network.
IV-E Computational Complexity
@@
In TABLE II, we compare the number of parameters of our framework, which includes $K$ modes and additional features, with the state-of-the-art ASTGCN method. It shows that the increase in parameters by adding features to the traffic flow is negligible. However, the time complexity in decoupling the traffic counts takes a significant amount of time relative to training the model. The complexity $\mathcal{O}$ in terms of time for each element of an ST block in ASTGCN is formulated in this section. Typically, $N\gg d$ and $N\gg T_{w}$ for targeted traffic prediction applications. The time complexity for normalized spatial attention $\mathbf{S}^{{}^{\prime}}$ is $\mathcal{O}(N^{3})$. The complexity for normalized temporal attention is $\mathcal{O}(NdT_{w})$, for Chebyshev convolution is $\mathcal{O}(T_{w}MN^{2}C)$, and for time convolution is $\mathcal{O}(FO_{h}O_{w}K_{h}K_{w}C)$, where
$O_{h}=\bigg{\lfloor}\frac{N-K_{h}}{S_{h}}\bigg{\rfloor}+1,$
$O_{w}=\bigg{\lfloor}\frac{T-K_{w}}{S_{w}}\bigg{\rfloor}+1,$
are the height and width of the output tensors of the 2D convolution, respectively. The width and height of the kernel are described by $K_{w}$ and $K_{h}$, and the strides of the convolution are described by $S_{w}$ and $S_{h}$. The number of the output filter $F$ of CNN and $C$ is the number of the Chebyshev channel. If we have the $B$ number of ST blocks in ASTGCN, the complexity will be $\mathcal{O}(B(N^{3}+NdT_{w}+T_{w}MN^{2}C+FO_{h}O_{w}K_{h}K_{w}C))$.
The computational complexity for the VMD includes: mirroring of the signal is $\mathcal{O}(L)$, the FFT $\mathcal{O}(L\log L)$, the update of the modes $\mathcal{O}(N_{\rm{iter}}KL)$, and the reconstruction of the signal is $\mathcal{O}(KL\log(L))$. Considering the number of iterations $N_{\rm{iter}}=500$, the overall complexity of VMD for all nodes is $\mathcal{O}(NN_{\rm{iter}}KL)$. It shows that the update of modes $u_{k}(\omega)$ and central frequencies $\omega_{k}$ quantify the computation cost of the decomposition of the signal.
Due to this limitation, our framework is well-suited for offline prediction of the ST data because we are decomposing the 1-year data rather than input samples. The computation of the decomposition of the signal can be optimized and made learnable by using deep unrolling of the VMD algorithm. This proposed approach can be implemented for future modifications of the framework.
V Conclusion
@@
In this work, we successfully implemented a hybrid method for both short-term and long-term spatio-temporal (ST) forecasting. The ST data is decomposed into modes, which are then used as features for the deep-learning model to predict future states. The number of modes for real-time applications is determined using the reconstruction loss. The low-frequency modes significantly contribute to horizon prediction while high-frequency modes contain the noise from the signal. By discarding the highest frequency mode from the features, we can enhance the prediction accuracy. However, the bandwidth and tolerance values significantly influence the convergence and performance of our proposed framework. We evaluated the performance of our architecture on the traffic flow dataset, it outperforms in both short and long-term prediction compared to state-of-the-art methods.

### v3

@@ 4.4. Ablative Study
The performance of the framework has been evaluated on various combinations of decomposition hyper-parameters such as $\alpha$, and $\epsilon$, for the SD region. TABLE 3 details different performance metrics for horizons $3$, $6$, and $12$, and the average. It has been observed that for $\alpha=1000$, the reconstruction loss is relatively lower compared to $\alpha=2000$. Changing the $\epsilon$ value does not significantly affect the reconstruction loss, indicating that the convergence is sensitive to initial conditions, which may lead to local minima instead of global minimum. For VMGCN-v2, it is noticed that decreasing the bandwidth constraint and increasing the tolerance value improves the performance metrics. The overall cumulative effect on the reconstruction loss term in the region remains unaffected by changes in the $\epsilon$ term. However, the difference in the reconstructed loss term of modes per node is defined as
(16) $\frac{1}{L}\lvert\sum_{i}u(t)_{i,\epsilon_{1}}-\sum_{i}u(t)_{i,\epsilon_{2}}\rvert,$
@@
where $L$ is the total length of the signal. For this comparison, we used $\epsilon_{1}=1\times 10^{-7}$ and $\epsilon_{2}=1\times 10^{-6}$. With a bandwidth constraint of $\alpha=1000$, the average loss term on the SD region is measured as $1.134\times 10^{-5}$, indicating a convergence difference between these sets of decomposition parameters, with different local minima being reached. The framework performs better with $\epsilon_{2}$ compared to $\epsilon_{1}$. If the bandwidth constraint is set at $\alpha=2000$, the $L_{1}$ loss term of the mode approaches zero. This indicates that the role of tolerance in convergence is less significant for a larger bandwidth value. From the table, we can also infer that the changing the bandwidth constraint has a dominant impact on the performance of the network. By eliminating the high-frequency (IMF13) in VMGCN-v1, the DNN is trained and the performance of the model is evaluated. It appears that the overall performance of the network is enhanced, specifically the long-term predictions. Furthermore, we observe that the high-frequency component adversely affects the long-term prediction capability of the network.
(a) Sensor ID: 1114091
(b) Sensor ID: 1108597
(c) Sensor ID: 1123256
Figure 11. This plot shows the comparison of ground truth with the predicted traffic flow for horizon 1 on three distinct locations in the SD region.
4.5. Computational Complexity
@@
In TABLE 2, we compare the number of parameters of our framework, which includes $K$ modes and additional features, with the state-of-the-art ASTGCN method. It shows that the increase in parameters by adding features to the traffic flow is negligible. However, the time complexity in decoupling the traffic counts takes a significant amount of time relative to training the model. The computational complexity for each element of an ST block in ASTGCN is formulated in this section. Typically, $N\gg d$ and $N\gg T_{w}$ for targeted traffic prediction applications. The time complexity for normalized spatial attention $\mathbf{S}^{{}^{\prime}}$ is $\mathcal{O}(N^{3})$. The complexity for normalized temporal attention is $\mathcal{O}(NdT_{w})$, for Chebyshev convolution is $\mathcal{O}(T_{w}MN^{2}d)$, and for time convolution is $\mathcal{O}(FO_{h}O_{w}K_{h}K_{w}C)$, where $O_{h}=\lfloor(N-K_{h})/S_{h}\rfloor+1,$ and $O_{w}=\lfloor({T-K_{w}})/{S_{w}}\rfloor+1$ are the height and width of the output tensors of the 2D convolution, respectively. The width and height of the kernel are described by $K_{w}$ and $K_{h}$, and the strides of the convolution are described by $S_{w}$ and $S_{h}$. The number of the CNN output filter is denoted by $F$ and $C$ is the number of the Chebyshev polynomial coefficients feature channels. If we have the $B$ number of ST blocks in ASTGCN, the complexity will be $\mathcal{O}(B(N^{3}+NdT_{w}+T_{w}MN^{2}d+FO_{h}O_{w}K_{h}K_{w}C))$.
The computational complexity for VMD involves several components when processing graph signals. The initial mirroring operation of the signal of length $L$ and the subsequent FFT require $\mathcal{O}(L)$ and $\mathcal{O}(L\log L)$ operations respectively. During the decomposition, updating each of the $K$ modes over $N_{\text{iter}}$ iterations incurs a cost of $\mathcal{O}(N_{\text{iter}}KL)$, while the final signal reconstruction through inverse FFTs requires $\mathcal{O}(KL\log L)$ operations. Combining these components, the per-node complexity simplifies to $\mathcal{O}(L\log L+N_{\text{iter}}KL)$. When extended to all $N$ graph nodes (processed independently) with the iteration bound $N_{\text{iter}}\leq N_{m}$, the overall VMD complexity becomes $\mathcal{O}(N(L\log L+N_{\text{iter}}KL))$. This formulation highlights that the $\mathcal{O}(NL\log L)$ term dominates for typical cases where $N_{\text{iter}}$ remains moderate relative to the signal length $L$. It also shows that the update of modes $\hat{u}_{k}(\omega)$ and central frequencies $\omega_{k}$ quantify the computation cost of the decomposition of the signal. The average CPU decomposition time for the SD, GBA, and GLA regions is 6 h, 27 h, and 102 h, respectively. Since we decompose the 1-year data rather than input samples, our framework is well-suited for offline prediction of the ST data. This work focuses on evaluating the dynamic modes and their impact on the accuracy of the model. In future work, we aim to explore asynchronous VMD to decompose batches of time series concurrently to reduce wall-clock decomposition time and make VMGCN suitable for real-time forecasting. The computation of the decomposition of the signal can also be optimized and made learnable by using deep unrolling of the VMD algorithm.
5. Conclusion
@@
The proposed VMGCN-driven forecasting is the first framework to integrate variational mode decomposition with attention-based graph convolutions, achieving both state-of-the-art accuracy and frequency-level interpretability in large-scale traffic forecasting. In this work, we have successfully implemented a hybrid method for both short-term and long-term spatiotemporal (ST) forecasting. The ST data is decomposed into modes, which are then used as features for the deep-learning model to predict future states. The number of modes for real-time applications is determined using the reconstruction loss. The low-frequency modes significantly contribute to horizon prediction while high-frequency modes contain the noise from the signal. By discarding the highest frequency mode from the features, we can enhance the prediction accuracy. However, the bandwidth and tolerance values significantly influence the convergence and performance of our proposed framework. We evaluated the performance of our architecture on the traffic flow dataset, it outperforms in both short and long-term prediction compared to state-of-the-art methods.